# Recent Advances and Trends in Research Paper Recommender Systems: A Comprehensive Survey

### v1

@@ 3.1.3 How do the RP-RSs represent the papers?
When analyzing the approaches implemented to represent the papers, it was observed that 11.11% of the works encode papers using graph-based strategies and 79.37% use Vector Space Model (VSM) representation techniques. Only one work represent the paper employing a text-based approach using keyphrases [53].
Regarding graph-based representation, only one work, [76], focuses solely on textual content, representing papers as nodes along with other content-related entities such as authors, keywords, venues, funds, and organizations. In contrast, several works combine textual content with relational information. For example, in [28, 29], Heterogeneous Information Networks (HINs) nodes include papers, authors, users, venues, and tags. [26] represents papers, authors, words, and topics as nodes, while [61] includes users, papers, terms, authors, and venues. Additionally, [80] defines the popularity network by weighting citation links according to the difference in popularity scores between cited and citing papers, emphasizing connections involving more popular articles.
@@
Works that rely solely on textual content predominantly employ transformer-like models. For instance, [62] leverage LLMs trained on scientific paper corpora, such as SciBERT, alongside BERT. In addition, [64] incorporate SciBERT and Recurrent Neural Networks (RNN). Lastly, [52] adopt SPECTER. However, some current works continue to apply classical embedding techniques such as Word2Vec [49] and Doc2Vec [16, 22], demonstrating their ongoing relevance alongside newer methods.
@@
Within this content-based embedding landscape, different techniques are combined to generate embeddings. [55] combine BERT with ResNets. [32] integrate RNNs and Convolution Neural Networks (CNNs). [56] explore fusing Word2Vec embeddings with RNNs.
@@
Works that combine textual content with relational information primarily rely on transformers-based models. Some use simpler approaches, employing only transformers to generate paper embeddings [17, 33]. In contrast, others integrate transformers with graph-based methods. For example, several works [18, 65, 40, 79, 54, 45, 30] combine transformers with Neural Network (NN)-based graph approaches, while [72] pairs transformers with meta-paths to generate final paper representations. Additionally, [71] integrates transformers with both NN-based graph and meta-path techniques. On the other hand, works that rely more heavily on meta-paths often combine them with various NN models. Specifically, [68] integrate meta-paths with RNNs. [73] combine meta-paths with Graph Neural Networks (GNNs). [77] combine meta-paths with both GNNs and CNNs. [78] combine meta-paths with RNNs and GNNs.
@@
17.46% of the works represent papers using Latent Feature Vectors (LFVs). A purely content-based approach is presented in [66], which uses TF-IDF. Works that combine textual content with relational information typically use Matrix Factorization (MF) methods. Some apply MF techniques alone, such as Alternating Least Squares (ALS), Bayesian Personalized Ranking (BPR), and Probabilistic Matrix Factorizatio (PMF) [25, 50]. However, most combine MF with text-based approaches to address the cold start problem. These include Topic Modeling (TM) [3, 24, 51] and transformers [60, 20]. Auto-encoders are also combined with MF [74, 57].
@@
Feature Vector (FV) and Topic Distribution (TD) represent 6.35% and 4.76% of the representations, respectively. For FV, TF-IDF is the main technique used in works that combine textual content with relational information [33, 19, 44]. Regarding TD, some works use solely content-based information by applying Latent Dirichlet Allocation (LDA) [13, 11], while others combine content-based and relational-based information [23].
@@
Finally, some works produce different paper representations using embeddings and FV techniques [38], which use purely content-based information, or FV and LFV techniques [50], which combine content-based and relational information, to create different recommendation sets that are then fused with hybridization techniques.
@@
Graph-based HIN - - HIN [61, 76, 29, 28, 26]
Factor analysis - HoIN(PN) [80]
HoIN - - HoIN(CN) [31]
VSM Embedding DL Transformer, GNN HoIN(CN) [30, 54]
HIN [65]
BG [18]
Transformer, GNN, Meta-path HIN [71]
Transformer, GCN HIN [40, 79]
HoIN(CN) [45]
Transformer, RNN, Meta-path, RW, TF-IDF HIN [72]
Transformer, RNN, CNN HIN(KG) [70]
Transformer, RNN - [64]
VSM Embedding DL Transformer, MLP HIN [4]
Transformer, Residual network - [55]
Transformer - [33, 52, 17, 42, 62]
NN - [22, 49, 16]
NN, RNN - [56]
NN, Topic modeling - [38]
GCN APRHG [81]
- [21]
CNN - [63]
GNN, CNN HoIN(CoN) [43]
GNN, Meta-path HIN [73]
HGNN, RW HIN [67]
RNN, Highway network, CNN - [32]
RNN, GNN, Meta-path HIN [78]
RNN, CNN, Meta-path HIN(AHG) [77]
RNN, Meta-path HIN [68]
RNN, GAT HIN [69]
DL,MF HAN, MF - [1]
VSM LFV DL,MF Transformer, Cosine similarity, MF Graph structure [20]
MLP, AutoEncoder, MF HIN(BG) [74]
Dual-AutoEncoder, MF - [57]
TF-IDF, GNN, MF HoIN(CN) [75]
Transformer, MF, CNN - [60]
MF Topic modeling, MF HoIN(CN) [24]
MF, Topic modeling - [3]
MF, TF-IDF - [66]
MF - [25, 50]
MF, Statistical method Topic modeling, MF HIN [51]
FV Weighted entity vector - - [50]
Statistical method TF-IDF - [33, 19, 44]
TF-IDF, Topic modeling - [38]
TF - [46]
Statistical method, Feature engineering Topic modeling - [15]
TD Statistical method Topic modeling - [23, 11, 13]
Text-based keyphrases - Keyphrase extraction - [53]
It is worth mentioning that a very small percentage of works use some kind of multi-representation approach, that is, the same paper has different representations (see Table LABEL:tab:PaperMultiRepresentation). VSM and graph-based node representations are combined to produce combined recommendations based in cosine similarity and citation similarity [58, 39].
@@
Vector Space Model Term frequency - - [58]
Graph-based HoIN(CN) - HoIN(CN) [58]
Vector Space Model Feature vector - - [39]
Graph-based HoIN(CN) - HoIN(CN) [39]
@@ 3.1.4 How do the RP-RSs represent the users?
61.9% of the analyzed works use some type of representation to model users that allows the generation of personalized recommendations. 10% encode users using graph-based strategies, 87.5% use VSM representation techniques and only one work used other kind of representation by generating a vectorizer and a classifier to define or represent the user [33].
@@
The subcategories and percentages of usage identified in the VSM are very similar to those for papers, except that, in this case, there are no TD representations as none of the works analyzed use this approach to model the profile of the users. 50% of the works represent users as embeddings, 30% use LFV to model users whilst 7.5% of the works use FV.
The works that use embedding representation techniques to model users rely mainly in graph-based techniques. NN-based approaches such as, Graph Convolution Networks (GCNs) [81, 21, 16], GNNs [18, 17], RNNs [45, 69], Multi Layer Perceptrons (MLPs) [70] or Random Walk (RW) [67]. [40, 71, 79] use GNNs along with transformers to get the user embeddings. Some works combine RNNs with meta-paths in order to model the users [68, 77, 72, 78].
However, approaches that do not rely on graph-based techniques can also be found among the works that rely on embedding representations, namely, Word2Vec [49], a combination of RNN with CNN [32] or transformers used with CNN [30].
There are works that only rely on MF to represent users [60, 1, 66, 25, 50]. Other works combine MF with TM [24, 51, 3], with auto-encoders [74, 57], transformers [20] or GNNs [75].
FV represent the 7.5% of the approaches, where TF-IDF [44] and TM [15] are the techniques used.
@@
Graph-based HIN - - HIN [61, 29, 28, 26]
VSM Embedding DL Transformer, CNN - [30]
Transformer, GNN, Meta-path HIN [71]
Transformer, GCN HIN [79, 40]
RNN, Meta-path HIN [68]
RNN, Meta-path, GNN HIN [78]
RNN, Meta-path, TF-IDF, RW HIN [72]
RNN, Highway network, CNN Graph structure [32]
RNN, GAT HIN [69]
RNN, GCN HoIN(CN) [45]
GCN APRHG [81]
- [21]
HIN(KG) [16]
GNN BG [18]
KG(HIN) [17]
GNN, MLP HIN(KG) [70]
HGNN, RW HIN [67]
VSM Embedding DL LSTM, CNN, Meta-path HIN(AHG) [77]
NN HIN [73]
- [49]
LFV MF MF - [1, 25, 50, 60, 66]
MF, Topic modeling HoIN(CN) [24]
MF, Topic modeling, ARM - [3]
MF, Transformer, Cosine similarity Graph type [20]
MF, MLP, AutoEncoder HIN(BG) [74]
MF, Dual-AutoEncoder - [57]
MF, GNN HoIN(CN) [75]
MF, Statistical method MF, Topic modeling HIN [51]
FV Statistical method, Feature engineering Topic modeling - [15]
VSM FV Statistical method TF-IDF - [44]
Weighted entity vector - - [50]
Other Vectorized & - - - [33]
@@ 3.1.5 How do the RP-RSs generate the recommendations?
All works that use the inner product-based method generate personalized recommendations by calculating the similarity between user and paper vector representations through the dot product. 52.17% of the works that use inner product are based on embedding representations using graph-based techniques as a part of a complex model. In these models, various training strategies are employed, ranging from well-established methods such as BPR loss [18, 67], cross-entropy loss [30, 40, 17, 68], and tuple-wise loss [81], to more specialized approaches [21, 16, 69]. Additionally, works utilizing transformer-based [55] and RNN-based embeddings [32], which adopt BPR loss and Weighted Mean Squared (WMS) loss respectively as training strategies, have also been identified.
@@
47.83% of the works that utilize inner products as part of the recommendation generation process involve complex models based on MF to produce user and paper representations. Some of these incorporate MF with auto-encoder architectures, either leveraging ALS to minimize the objective function [74] or using the cross-entropy loss [57]. Others combine transformers with MF models, applying BPR loss [20] or Mean Square Error (MSE) loss [60]. In certain cases, when recommendations are generated for a group of users, the Evidential Reasoning (ER) rule is applied to aggregate member ratings [66] before computing the inner product. Finally, works that incorporate GNN and Hierarchical Attention Network (HAN) are integrated with MF and inner products, with cross-entropy loss [75] applied in the former and Weighted Approximate-Rank Pairwise (WARP) loss [1] in the latter.
@@
19.05% of the works utilize a ML approach, 66.6% of them producing personalized recommendations. Most of them employ a MLP framework, with several incorporating cross-entropy loss [71, 78], while others opt for ad-hoc loss functions [70, 77]. In addition to MLP, a work combining RNN and Support Vector Machine (SVM) models [15], a GCN integrated with BPR loss for training [73], and a NN with cross-entropy loss [79] have also been found to yield personalized recommendations. Those works that do not provide personalized recommendations (33.3%) typically employ various NN architectures. [65] uses cross-entropy loss, while [64] uses MSE loss, and [54] adopts a customized loss function.
@@
28.57% of the works rely on similarity-based methods, 77.78% of them producing personalized recommendations. The most used is cosine similarity [52, 43, 22, 46, 38, 11, 62, 49]. [39, 58] use both cosine similarity and citation similarity. [53] applies cosine similarity along with Jaccard similarity. [72] also uses cosine similarity after applying a clustering process to reduce the computation time. [63] uses contrastive loss with cosine similarity to increase similarity within document paragraphs and decrease similarity across different documents. In [42] the hierarchical loss balances global themes and fine-grained aspects by jointly optimizing general and aspect-specific embeddings. Other similarity measures have also been identified in the analysis. [76] combines PathSim with RW. [13] uses Jensen-Shannon Divergence (JSD), while [23] integrates JSD with PageRank.
@@
Works employing graph-based approaches to generate recommendations utilize RW algorithms with restart [28, 29, 26] or without restart[61]. All of them provide personalized recommendations. [56] is the only graph-based approach which does not provide personalized recommendations. This work uses the COOT optimizing algorithm to efficiently identify highly influential papers.
@@
Finally, some other approaches, which represent 7.94% of the total analyzed works, have been identified. Only one of these works presents a personalized recommendation system. This work, [50], uses inner product and a relevance score to generate the recommendations. The rest of the works found in this group use co-citation and bibliographic coupling [31]. The hybrid solution presented in [19] combines bibliographic coupling, cosine similarity, PubMed Related Article (PMRA), and BM25. Other approaches use a personalized score [80] or a Naïve Bayes classifier [33].

### v2

@@ 3.1.3 How do the RP-RSs represent the papers?
When analyzing the approaches implemented to represent the papers, it was observed that 11.11% of the works encode papers using graph-based strategies and 79.37% use Vector Space Model (VSM) representation techniques. Only one work represent the paper employing a text-based approach using keyphrases [IJ:IEEE:SARWAR:2021].
Regarding graph-based representation, only one work, [IJ:AI:Xiao:2023], focuses solely on textual content, representing papers as nodes along with other content-related entities such as authors, keywords, venues, funds, and organizations. In contrast, several works combine textual content with relational information. For example, in [IC:IKT:JAFARI:2021, IJ:AUT:Jafari:2023], Heterogeneous Information Networks (HINs) nodes include papers, authors, users, venues, and tags. [IJ:Scientometrics:Huang:2024] represents papers, authors, words, and topics as nodes, while [PP:Mahesh:2023] includes users, papers, terms, authors, and venues. Additionally, [IJ:JOI:Yadav:2022] defines the popularity network by weighting citation links according to the difference in popularity scores between cited and citing papers, emphasizing connections involving more popular articles.
@@
Works that rely solely on textual content predominantly employ transformer-like models. For instance, [EC:ECIR:Takahashi:2022] leverage LLMs trained on scientific paper corpora, such as SciBERT, alongside BERT. In addition, [IJ:Scientometrics:Thierry:2023] incorporate SciBERT and Recurrent Neural Networks (RNN). Lastly, [IC:WIST:Roßner:2023] adopt SPECTER. However, some current works continue to apply classical embedding techniques such as Word2Vec [AC:ACIIDS:PAN:2024] and Doc2Vec [IJ:AS:Chen:2023, IJ:Scientometrics:GUNDOGAN:2022], demonstrating their ongoing relevance alongside newer methods.
@@
Within this content-based embedding landscape, different techniques are combined to generate embeddings. [IC:ASSP:Shen:2024] combine BERT with ResNets. [arXiv:KARIMI:2022] integrate RNNs and Convolution Neural Networks (CNNs). [IJ:DSM:Sivasankari:2023] explore fusing Word2Vec embeddings with RNNs.
@@
Works that combine textual content with relational information primarily rely on transformers-based models. Some use simpler approaches, employing only transformers to generate paper embeddings [IJ:AI:GAO:2023, IJ:DATABASE:Kart:2022]. In contrast, others integrate transformers with graph-based methods. For example, several works [IC:MLBDBI:GUAN:2022, IJ:AI:Thierry:2023, CF:CSCW:LI:2022, IC:ICDE:XIE:2022, ARXIV:Shen:2024, IC:EITCE:NIU:2023, IJ:Scientometrics:Jiang:2023] combine transformers with Neural Network (NN)-based graph approaches, while [IJ:JES:LI:2024] pairs transformers with meta-paths to generate final paper representations. Additionally, [IC:SEAI:Wang:2024] integrates transformers with both NN-based graph and meta-path techniques. On the other hand, works that rely more heavily on meta-paths often combine them with various NN models. Specifically, [IC:DASFAA:Wang:2022] integrate meta-paths with RNNs. [IS:CSET:WU:2022] combine meta-paths with Graph Neural Networks (GNNs). [IJ:ESWA:XIAO:2023] combine meta-paths with both GNNs and CNNs. [IJ:KBS:Xiao:2023] combine meta-paths with RNNs and GNNs.
@@
17.46% of the works represent papers using Latent Feature Vectors (LFVs). A purely content-based approach is presented in [IJ:SMC:WANG:2022], which uses TF-IDF. Works that combine textual content with relational information typically use Matrix Factorization (MF) methods. Some apply MF techniques alone, such as Alternating Least Squares (ALS), Bayesian Personalized Ranking (BPR), and Probabilistic Matrix Factorizatio (PMF) [J:EIJ:Huang:2024, IJ:IS:PINEDO:2024]. However, most combine MF with text-based approaches to address the cold start problem. These include Topic Modeling (TM) [IJ:CDS:AHMEDI:2022, IJ:JS:Huang:2023, IJ:IEEE:REN:2022] and transformers [IC:ICITACEE:Switrayana:2022, IJ:TIIS:GUO:2024]. Auto-encoders are also combined with MF [IJ:ESWA:XI:2023, IJ:KIS:STERGIOPOULOS:2024].
@@
Feature Vector (FV) and Topic Distribution (TD) represent 6.35% and 4.76% of the representations, respectively. For FV, TF-IDF is the main technique used in works that combine textual content with relational information [IJ:DATABASE:Kart:2022, IJ:SHTI:GUO:2022, J:JIS:MUNGEN:2022]. Regarding TD, some works use solely content-based information by applying Latent Dirichlet Allocation (LDA) [IJ:EIT:Smail:2023, IJ:ELECTRONICS:Blazevic:2023], while others combine content-based and relational-based information [IJ:IEEE:HADHIATMA:2023].
@@
Finally, some works produce different paper representations using embeddings and FV techniques [IJ:AMEC:KUS:2023], which use purely content-based information, or FV and LFV techniques [IJ:IS:PINEDO:2024], which combine content-based and relational information, to create different recommendation sets that are then fused with hybridization techniques.
@@
Graph-based HIN - - HIN [PP:Mahesh:2023, IJ:AI:Xiao:2023, IJ:AUT:Jafari:2023, IC:IKT:JAFARI:2021, IJ:Scientometrics:Huang:2024]
Factor analysis - HoIN(PN) [IJ:JOI:Yadav:2022]
HoIN - - HoIN(CN) [IJ:Scientometrics:Kanwal:2024]
VSM Embedding DL Transformer, GNN HoIN(CN) [IJ:Scientometrics:Jiang:2023, ARXIV:Shen:2024]
HIN [IJ:AI:Thierry:2023]
BG [IC:MLBDBI:GUAN:2022]
Transformer, GNN, Meta-path HIN [IC:SEAI:Wang:2024]
Transformer, GCN HIN [CF:CSCW:LI:2022, IC:ICDE:XIE:2022]
HoIN(CN) [IC:EITCE:NIU:2023]
Transformer, RNN, Meta-path, RW, TF-IDF HIN [IJ:JES:LI:2024]
Transformer, RNN, CNN HIN(KG) [IJ:AS:Wang:2024]
Transformer, RNN - [IJ:Scientometrics:Thierry:2023]
VSM Embedding DL Transformer, MLP HIN [IJ:Scientometrics:ALI:2022]
Transformer, Residual network - [IC:ASSP:Shen:2024]
Transformer - [IJ:DATABASE:Kart:2022, IC:WIST:Roßner:2023, IJ:AI:GAO:2023, IJ:IPM:Long:2024, EC:ECIR:Takahashi:2022]
NN - [IJ:Scientometrics:GUNDOGAN:2022, AC:ACIIDS:PAN:2024, IJ:AS:Chen:2023]
NN, RNN - [IJ:DSM:Sivasankari:2023]
NN, Topic modeling - [IJ:AMEC:KUS:2023]
GCN APRHG [IC:KSEM:YU:2022]
- [IJ:mathematics:GUO:2024]
CNN - [IJ:Scientometrics:TAN:2023]
GNN, CNN HoIN(CoN) [IJ:ESWA:MEI:2022]
GNN, Meta-path HIN [IS:CSET:WU:2022]
HGNN, RW HIN [IJ:ASC:Wang:2024]
RNN, Highway network, CNN - [arXiv:KARIMI:2022]
RNN, GNN, Meta-path HIN [IJ:KBS:Xiao:2023]
RNN, CNN, Meta-path HIN(AHG) [IJ:ESWA:XIAO:2023]
RNN, Meta-path HIN [IC:DASFAA:Wang:2022]
RNN, GAT HIN [IJ:ESWA:Wang:2023]
DL,MF HAN, MF - [ARXIV:Mohamed:2022]
VSM LFV DL,MF Transformer, Cosine similarity, MF Graph structure [IJ:TIIS:GUO:2024]
MLP, AutoEncoder, MF HIN(BG) [IJ:ESWA:XI:2023]
Dual-AutoEncoder, MF - [IJ:KIS:STERGIOPOULOS:2024]
TF-IDF, GNN, MF HoIN(CN) [IJ:SC:XI:2024]
Transformer, MF, CNN - [IC:ICITACEE:Switrayana:2022]
MF Topic modeling, MF HoIN(CN) [IJ:JS:Huang:2023]
MF, Topic modeling - [IJ:CDS:AHMEDI:2022]
MF, TF-IDF - [IJ:SMC:WANG:2022]
MF - [J:EIJ:Huang:2024, IJ:IS:PINEDO:2024]
MF, Statistical method Topic modeling, MF HIN [IJ:IEEE:REN:2022]
FV Weighted entity vector - - [IJ:IS:PINEDO:2024]
Statistical method TF-IDF - [IJ:DATABASE:Kart:2022, IJ:SHTI:GUO:2022, J:JIS:MUNGEN:2022]
TF-IDF, Topic modeling - [IJ:AMEC:KUS:2023]
TF - [IJ:SECS:Hamisu:2024]
Statistical method, Feature engineering Topic modeling - [IJ:IS:Chaudhuri:2022]
TD Statistical method Topic modeling - [IJ:IEEE:HADHIATMA:2023, IJ:ELECTRONICS:Blazevic:2023, IJ:EIT:Smail:2023]
Text-based keyphrases - Keyphrase extraction - [IJ:IEEE:SARWAR:2021]
It is worth mentioning that a very small percentage of works use some kind of multi-representation approach, that is, the same paper has different representations (see Table LABEL:tab:PaperMultiRepresentation). VSM and graph-based node representations are combined to produce combined recommendations based in cosine similarity and citation similarity [IJ:IEEE:STERLING:2022, IJ:Heliyon:LI:2024].
@@
Vector Space Model Term frequency - - [IJ:IEEE:STERLING:2022]
Graph-based HoIN(CN) - HoIN(CN) [IJ:IEEE:STERLING:2022]
Vector Space Model Feature vector - - [IJ:Heliyon:LI:2024]
Graph-based HoIN(CN) - HoIN(CN) [IJ:Heliyon:LI:2024]
@@ 3.1.4 How do the RP-RSs represent the users?
61.9% of the analyzed works use some type of representation to model users that allows the generation of personalized recommendations. 10% encode users using graph-based strategies, 87.5% use VSM representation techniques and only one work used other kind of representation by generating a vectorizer and a classifier to define or represent the user [IJ:DATABASE:Kart:2022].
@@
The subcategories and percentages of usage identified in the VSM are very similar to those for papers, except that, in this case, there are no TD representations as none of the works analyzed use this approach to model the profile of the users. 50% of the works represent users as embeddings, 30% use LFV to model users while 7.5% of the works use FV.
The works that use embedding representation techniques to model users rely mainly in graph-based techniques. NN-based approaches such as, Graph Convolution Networks (GCNs) [IC:KSEM:YU:2022, IJ:mathematics:GUO:2024, IJ:AS:Chen:2023], GNNs [IC:MLBDBI:GUAN:2022, IJ:AI:GAO:2023], RNNs [IC:EITCE:NIU:2023, IJ:ESWA:Wang:2023], Multi Layer Perceptrons (MLPs) [IJ:AS:Wang:2024] or Random Walk (RW) [IJ:ASC:Wang:2024]. [CF:CSCW:LI:2022, IC:SEAI:Wang:2024, IC:ICDE:XIE:2022] use GNNs along with transformers to get the user embeddings. Some works combine RNNs with meta-paths in order to model the users [IC:DASFAA:Wang:2022, IJ:ESWA:XIAO:2023, IJ:JES:LI:2024, IJ:KBS:Xiao:2023].
However, approaches that do not rely on graph-based techniques can also be found among the works that rely on embedding representations, namely, Word2Vec [AC:ACIIDS:PAN:2024], a combination of RNN with CNN [arXiv:KARIMI:2022] or transformers used with CNN [IJ:Scientometrics:Jiang:2023].
There are works that only rely on MF to represent users [IC:ICITACEE:Switrayana:2022, ARXIV:Mohamed:2022, IJ:SMC:WANG:2022, J:EIJ:Huang:2024, IJ:IS:PINEDO:2024]. Other works combine MF with TM [IJ:JS:Huang:2023, IJ:IEEE:REN:2022, IJ:CDS:AHMEDI:2022], with auto-encoders [IJ:ESWA:XI:2023, IJ:KIS:STERGIOPOULOS:2024], transformers [IJ:TIIS:GUO:2024] or GNNs [IJ:SC:XI:2024].
FV represent the 7.5% of the approaches, where TF-IDF [J:JIS:MUNGEN:2022] and TM [IJ:IS:Chaudhuri:2022] are the techniques used.
@@
Graph-based HIN - - HIN [PP:Mahesh:2023, IJ:AUT:Jafari:2023, IC:IKT:JAFARI:2021, IJ:Scientometrics:Huang:2024]
VSM Embedding DL Transformer, CNN - [IJ:Scientometrics:Jiang:2023]
Transformer, GNN, Meta-path HIN [IC:SEAI:Wang:2024]
Transformer, GCN HIN [IC:ICDE:XIE:2022, CF:CSCW:LI:2022]
RNN, Meta-path HIN [IC:DASFAA:Wang:2022]
RNN, Meta-path, GNN HIN [IJ:KBS:Xiao:2023]
RNN, Meta-path, TF-IDF, RW HIN [IJ:JES:LI:2024]
RNN, Highway network, CNN Graph structure [arXiv:KARIMI:2022]
RNN, GAT HIN [IJ:ESWA:Wang:2023]
RNN, GCN HoIN(CN) [IC:EITCE:NIU:2023]
GCN APRHG [IC:KSEM:YU:2022]
- [IJ:mathematics:GUO:2024]
HIN(KG) [IJ:AS:Chen:2023]
GNN BG [IC:MLBDBI:GUAN:2022]
KG(HIN) [IJ:AI:GAO:2023]
GNN, MLP HIN(KG) [IJ:AS:Wang:2024]
HGNN, RW HIN [IJ:ASC:Wang:2024]
VSM Embedding DL LSTM, CNN, Meta-path HIN(AHG) [IJ:ESWA:XIAO:2023]
NN HIN [IS:CSET:WU:2022]
- [AC:ACIIDS:PAN:2024]
LFV MF MF - [ARXIV:Mohamed:2022, J:EIJ:Huang:2024, IJ:IS:PINEDO:2024, IC:ICITACEE:Switrayana:2022, IJ:SMC:WANG:2022]
MF, Topic modeling HoIN(CN) [IJ:JS:Huang:2023]
MF, Topic modeling, ARM - [IJ:CDS:AHMEDI:2022]
MF, Transformer, Cosine similarity Graph type [IJ:TIIS:GUO:2024]
MF, MLP, AutoEncoder HIN(BG) [IJ:ESWA:XI:2023]
MF, Dual-AutoEncoder - [IJ:KIS:STERGIOPOULOS:2024]
MF, GNN HoIN(CN) [IJ:SC:XI:2024]
MF, Statistical method MF, Topic modeling HIN [IJ:IEEE:REN:2022]
FV Statistical method, Feature engineering Topic modeling - [IJ:IS:Chaudhuri:2022]
VSM FV Statistical method TF-IDF - [J:JIS:MUNGEN:2022]
Weighted entity vector - - [IJ:IS:PINEDO:2024]
Other Vectorized & - - - [IJ:DATABASE:Kart:2022]
@@ 3.1.5 How do the RP-RSs generate the recommendations?
All works that use the inner product-based method generate personalized recommendations by calculating the similarity between user and paper vector representations through the dot product. 52.17% of the works that use inner product are based on embedding representations using graph-based techniques as a part of a complex model. In these models, various training strategies are employed, ranging from well-established methods such as BPR loss [IC:MLBDBI:GUAN:2022, IJ:ASC:Wang:2024], cross-entropy loss [IJ:Scientometrics:Jiang:2023, CF:CSCW:LI:2022, IJ:AI:GAO:2023, IC:DASFAA:Wang:2022], and tuple-wise loss [IC:KSEM:YU:2022], to more specialized approaches [IJ:mathematics:GUO:2024, IJ:AS:Chen:2023, IJ:ESWA:Wang:2023]. Additionally, works utilizing transformer-based [IC:ASSP:Shen:2024] and RNN-based embeddings [arXiv:KARIMI:2022], which adopt BPR loss and Weighted Mean Squared (WMS) loss respectively as training strategies, have also been identified.
@@
47.83% of the works that utilize inner products as part of the recommendation generation process involve complex models based on MF to produce user and paper representations. Some of these incorporate MF with auto-encoder architectures, either leveraging ALS to minimize the objective function [IJ:ESWA:XI:2023] or using the cross-entropy loss [IJ:KIS:STERGIOPOULOS:2024]. Others combine transformers with MF models, applying BPR loss [IJ:TIIS:GUO:2024] or Mean Square Error (MSE) loss [IC:ICITACEE:Switrayana:2022]. In certain cases, when recommendations are generated for a group of users, the Evidential Reasoning (ER) rule is applied to aggregate member ratings [IJ:SMC:WANG:2022] before computing the inner product. Finally, works that incorporate GNN and Hierarchical Attention Network (HAN) are integrated with MF and inner products, with cross-entropy loss [IJ:SC:XI:2024] applied in the former and Weighted Approximate-Rank Pairwise (WARP) loss [ARXIV:Mohamed:2022] in the latter.
@@
19.05% of the works utilize a ML approach, 66.6% of them producing personalized recommendations. Most of them employ a MLP framework, with several incorporating cross-entropy loss [IC:SEAI:Wang:2024, IJ:KBS:Xiao:2023], while others opt for ad-hoc loss functions [IJ:AS:Wang:2024, IJ:ESWA:XIAO:2023]. In addition to MLP, a work combining RNN and Support Vector Machine (SVM) models [IJ:IS:Chaudhuri:2022], a GCN integrated with BPR loss for training [IS:CSET:WU:2022], and a NN with cross-entropy loss [IC:ICDE:XIE:2022] have also been found to yield personalized recommendations. Those works that do not provide personalized recommendations (33.3%) typically employ various NN architectures. [IJ:AI:Thierry:2023] uses cross-entropy loss, while [IJ:Scientometrics:Thierry:2023] uses MSE loss, and [ARXIV:Shen:2024] adopts a customized loss function.
@@
28.57% of the works rely on similarity-based methods, 77.78% of them producing personalized recommendations. The most used is cosine similarity [IC:WIST:Roßner:2023, IJ:ESWA:MEI:2022, IJ:Scientometrics:GUNDOGAN:2022, IJ:SECS:Hamisu:2024, IJ:AMEC:KUS:2023, IJ:ELECTRONICS:Blazevic:2023, EC:ECIR:Takahashi:2022, AC:ACIIDS:PAN:2024]. [IJ:Heliyon:LI:2024, IJ:IEEE:STERLING:2022] use both cosine similarity and citation similarity. [IJ:IEEE:SARWAR:2021] applies cosine similarity along with Jaccard similarity. [IJ:JES:LI:2024] also uses cosine similarity after applying a clustering process to reduce the computation time. [IJ:Scientometrics:TAN:2023] uses contrastive loss with cosine similarity to increase similarity within document paragraphs and decrease similarity across different documents. In [IJ:IPM:Long:2024] the hierarchical loss balances global themes and fine-grained aspects by jointly optimizing general and aspect-specific embeddings. Other similarity measures have also been identified in the analysis. [IJ:AI:Xiao:2023] combines PathSim with RW. [IJ:EIT:Smail:2023] uses Jensen-Shannon Divergence (JSD), while [IJ:IEEE:HADHIATMA:2023] integrates JSD with PageRank.
@@
Works employing graph-based approaches to generate recommendations utilize RW algorithms with restart [IC:IKT:JAFARI:2021, IJ:AUT:Jafari:2023, IJ:Scientometrics:Huang:2024] or without restart[PP:Mahesh:2023]. All of them provide personalized recommendations. [IJ:DSM:Sivasankari:2023] is the only graph-based approach which does not provide personalized recommendations. This work uses the COOT optimizing algorithm to efficiently identify highly influential papers.
@@
Finally, some other approaches, which represent 7.94% of the total analyzed works, have been identified. Only one of these works presents a personalized recommendation system. This work, [IJ:IS:PINEDO:2024], uses inner product and a relevance score to generate the recommendations. The rest of the works found in this group use co-citation and bibliographic coupling [IJ:Scientometrics:Kanwal:2024]. The hybrid solution presented in [IJ:SHTI:GUO:2022] combines bibliographic coupling, cosine similarity, PubMed Related Article (PMRA), and BM25. Other approaches use a personalized score [IJ:JOI:Yadav:2022] or a Naïve Bayes classifier [IJ:DATABASE:Kart:2022].